OpenStudy (anonymous):

what is the solution of l2x-1l=1

OpenStudy (anonymous):

In General |y|=a means (y=+a & y=-a) here y=2x-1 & a=1 It can be rewritten as =>2x-1=1 & 2x-1=-1 =>x=1 & x=0 are the solutions

OpenStudy (jdoe0001):

|2x-1| = 1 will have 2 solutions as Anu2401 already showed one solution will be by solving +1(2x-1) = 1 other solution will be by solving -(2x-1) = 1

OpenStudy (jdoe0001):

or -1(2x-1) = 1

OpenStudy (anonymous):

thank you